Sol Summary Report – January 26th

Crew 187 Sol 14 Summary Report 26JAN2018

Sol 14

Summary Title: In the end

Author’s name: Cynthia Fuertes Panizo

Mission Status: All nominal

Sol Activity Summary:

Hello Earth,

How is everything going on there?

Stay on Mars, it’s a dream that every Martian has. It is funny and enjoyable to stay here but also tired. Today we woke up in different hours. By 09:30, every Martian was awake. We ate breakfast; then, Oscar, Danton, Tania and I were getting ready for the last adventure as Martians. At 10:35, we started our last exploration; this time we found the right way at once, we arrived at the unknown canyon at 11:45. Again… we are the first crew in there… That feeling is awesome!!! You can’t imagine that kind of sensation. We would like to give it a name. Despacito road and Porg canyon, because, we consider that the road is dangerous so you have to drive slowly (slowly in Spanish is Despacito). In addition, It’s important to mention that Porg is the official pet of the crew 187, he always stays with us in all the adventures; therefore, we would like to leave a mark of our official pet.

We strongly recommend you to go to Porg Canyon! And drive very slowly in Despacito road, it’s for your security! The view in Porg canyon is beautiful, it’s a geological rich area. Thanks Shannon for recommending us this place!

At 14:45, we were inside the Hab. One hour and a half later, well, exactly talking, at 16:10, the extraordinary EVA began, at 17:10, they came back to home.

Later, Danton fly the drone around the Hab to do the aerial 3D mapping. At 18:20, wow, the moment that none Martian wants to arrive, arrived!… From that time, we were out of sim! But we feel like we are still Martians, at least, we are going to feel it until we left the MDRS and come back to Earth.

Currently, every Martian is running doing reports and the Mission Summary. Later we are going to have our last dinner together as crew 187. Tomorrow we are going to receive the next crew… Crew 188, we hope that you enjoy staying on Mars as much as we did!

Ad Astra,

Cynthia Fuertes Panizo

Commander Crew 187 – MDRS

Look Ahead Plan:

Receive the next crew… Crew 188, you have less than 24 hours on Earth!… Tomorrow you will be on Mars! This place is amazing, and you can learn a lot of things from the MDRS, every Martian, the landscape, and so on!

Furthermore, we are planning to enjoy our last day on Mars!

Anomalies in work:

None

Weather: Temperature: -5°C, Fair-weather, Winds speed: ESE 8 KPH, Humidity 35%, Barometer 30.29 in

Crew Physical Status: Healthy

Today’s EVA:

Location: Porg Canyon

Duration: From 10:35 until 14:45

Purpose: Testing of long-range EVA protocols and testing the Guache rover.

Today’s Extraordinary EVA:

Location: Quary Rd

Duration: From 16:10 until 17:10

Purpose: Usage of rovers for battery discharge.

Reports to be filed: Sol Summary, GeenHab Report, Operations Report, EVA Request, EVA Report, Astronomy Report, Journalist Report and Mission Summary.

Support Requested:

None

Greenhab Report – January 26th

GreenHab Report

Hernán David Mateus Jiménez

26Jan2018

Environmental control:

Heating

Ambient with door opening

Shade cloth on

Working Hour: 19:00
Inside temp at working hour: 15° C
Outside temp during working hours: 1° C
Inside temperature high: 29° C
Inside temperature low: 14° C
Inside humidity: 49 %RH

Inside humidity high: 49 %RH
Inside humidity low: 16 %RH

Hours of supplemental light:

For the crops 05:00 to 11:59 PM

Changes to crops: none

Daily water usage for crops: 8 gallons

Time(s) of watering for crops: 18:15

Changes to research plants: The plans of quinoa over the lysimeter are withered.

Aquaponics: Atila is assembling it. He is waiting for the amount of water to fill the tank, this is necessary to do the test.

Narrative: The last day simulation, I started disassembling the Lysimeter, Atila and I replanted 4 tomato plants in bigger pots and started to clean the greenhab.

In the afternoon, I realized that the first digit of the temperature on the screen of the sensor is not showing.

EVA Report – January 26th

EVA Report for 26Jan2018

Author: Oscar Ojeda

Start time: 10:00

End time: 15:00

Purpose of EVA: Aerial mapping of terrain, testing of long-range EVA protocols, testing of temperature regulation vest, Collection of possible halophile samples, Testing of the Guache Rover.

Destination: Unnamed zone on 1101 marker, west of Cow Dung Road, near white rock canyon

Coordinates: 519000 E, 4247500 N

Participants: Danton, Oscar, Tania, and Cynthia

Narrative: Today we were going to a place that was never explored by a crew before, and we took out our official pet, the Porg, with us. The crew left the airlock at 10:35 and headed south in 4 ATVs, upon finding the pole 1101 trough Cow Dung road, after about half an hour of traverse, we proceeded to drive into a road that hasn’t been used in a long while. The Road was very harsh, and we had to go very slow, so we decided to name the road “Despacito Rd”. Upon arriving we found several areas to explore, each one as beautiful as the other. Then is was when we realized that it was our last exploration EVA, and our goodbye from Mars. We walked through a canyon, which we decided to name Porg Canyon. The area is very rich geologically, with several sedimentary deposits, as well as clast rocks. There was also evidence of herbivore animals. The crew walked for 2 hours in the canyon and returned to the ATVs, starting the road home. Midway, the crew stopped in a rocky plain to test the functionality of the Guache Rover, using a wheel that was 3D printed in metal by the company ITAMCO, and designed with Purdue University. After that the crew resumed the way home, entering the airlock at 14:45.

No incident or anomaly was sustained during the EVA.

Operations Report – January 26th

Crew 187 Operations Report 26012018

SOL: 14

Name of person filing report: Luis Jose Antonio Diaz Lopez

Non-nominal systems: None.

Notes on non-nominal systems: Generator system limping along with a now-nominal routine.

Generator (hours run): 14.0h

Generator turned off, charging battery at 09h:50m

Generator turned on at 20h:20m3

Solar— SOC

@ 09:43: 54%

@ 19:43: 32%

Diesel: 94%

Propane: 88%

Ethanol Free Gasoline (5 Gallon containers for ATV): 10 Gallons

Water (trailer): 0 Gallons

Water (static): 60 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Water Meter: 01303618 Gallons

Toilet tank emptied: No

ATVs Used: HONDA, 350(1), 350(2), 350(3)

Oil Added? No

ATV Fuel Used: 7%

# Hours the ATVs were used today: 01h 40min

Notes on ATVs: None

Deimos rover used: Taked by the director to testing due the recently battery problems.

Hours: 102.3 (crew) + Director discretional hours

Beginning charge: 100%

Ending charge: –

Currently charging: Maybe

Sojourner rover used: Assigned to director only.

Hours: Director discretional hours

Beginning charge:

Ending charge: –

Currently charging: Maybe

Spirit rover used: Yes

Hours: 15.1h

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

Opportunity rover used: Yes

Hours: 12.4h

Beginning charge: 100%

Ending charge:

Currently charging: No

Curiosity rover used: Yes

Hours: 6.6h

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

HabCar used and why, where? No

General notes and comments: None

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of EVA operations: All nominal

Summary of GreenHab operations: All nominal

Summary of Science Dome operations: All nominal.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: Request: More water for the next crew.

Operations Report – January 25th

Operations Report

Subject Line: Crew 187 Operations Report 25JAN2018

Crew 187 Operations Report 25012018

SOL: 13

Name of person filing report: Luis Jose Antonio Diaz Lopez

Non-nominal systems: None.

Notes on non-nominal systems: Generator system limping along with a now-nominal routine.

Generator (hours run): 15.8h

Generator turned off, charging battery at 10h:40m

Generator turned on at 19h:43m

Solar— SOC

@ 10:10: 68%

@ 19:43: 37%

Diesel: 96%

Propane: 89%

Ethanol Free Gasoline (5 Gallon containers for ATV): 10 Gallons

Water (trailer): 0 Gallons

Water (static): 140 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Water Meter: 01302996 Gallons

Toilet tank emptied: No

ATVs Used: HONDA, 350(1), 350(2), 350(3)

Oil Added? No

ATV Fuel Used: 6%

# Hours the ATVs were used today: 01h 35min

Notes on ATVs: None

Deimos rover used: Taked by the director to testing due the recently battery problems.

Hours: 102.3 h

Beginning charge: 100%

Ending charge: –

Currently charging: Maybe

Sojourner rover used: Assigned to director only.

Hours: Director discretional hours

Beginning charge:

Ending charge: –

Currently charging: Maybe

Spirit rover used: Yes

Hours: 14.5h

Beginning charge: 100%

Ending charge: –

Currently charging: No

Opportunity rover used: Yes

Hours: 11.7h

Beginning charge: 100%

Ending charge:

Currently charging: Yes

Curiosity rover used: Yes

Hours: 5.9h

Beginning charge: 100%

Ending charge: –

Currently charging: Yes

HabCar used and why, where? No

General notes and comments: Opportunity have more hours than the other rovers because it was left turned on by error during the before EVA. We got 10 gallons of gasoline as backup for ATV’s.

Summary of internet: All nominal

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of EVA operations: All nominal

Summary of GreenHab operations: All nominal

Summary of Science Dome operations: All nominal.

Summary of RAM operations: Not Operational. Some generator supplies were delivered into the RAM.

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: None

Sol Summary – January 25th

Crew 187 Sol 13 Summary Report 25JAN2018

Sol 13

Summary Title: Breaking the habits

Author’s name: Cynthia Fuertes Panizo

Mission Status: All nominal

Sol Activity Summary:

Hi Earth,

How is everything going on there?

Last night we went to see the starry sky of Mars; however, the cloudy sky did not allow us to observe the beautiful Martian sky. There are only a few days left on Mars, if it depends on us, we would stay on Mars for much longer. The fatigue is felt in the Martians; on Mars, the hours pass faster than Earth. We slept late talking, coordinating things, moving forward in our projects and so on. Today we woke up at different times, but at 9:30 we were all ready to take breakfasts. One hour later, the Martian explorers, Atila, Luis, Tania and David started their adventure. Our explorers arrived at Mountain Goat Trail when they realized that they were in the wrong way to Candor Chasma. Then, they came back almost to Galileo Road 104… OMG!… again! The Martian explorers, couldn’t find the right way, as yesterday, the third one is the defeated. Therefore, at 12:03, they arrived at the right point and walked for 100 minutes until they came back to the rovers.

Meanwhile, in the Hab, Danton, Oscar and I got ready for the gala dinner. Of course, the chef in charge of the kitchen was our Martian chef Oscar. Danton and I were in charge of the cleaning. When the Martian explorers arrived at the Hab, we ate lunch together. Later, David helped me taking pictures for my project. At 15:55, Tania, Luis and I were in the airlock ready for the Extraordinary EVA, we arrived until the intersection between White Moon and Factory Butte View before we came back to our Martian Home.

Currently, we are doing the reports for Capcom and getting ready for our gala dinner. Today our special guest is Dr. Ruppert, we hope she enjoys her visit in our home.

In 2 hours, we are planning to use the Celestron Telescope to see the Martian Sky. Unfortunately, yesterday we couldn’t see anything at night, the sky was cloudy.

Ad Astra,

Cynthia Fuertes Panizo

Commander Crew 187 – MDRS

Look Ahead Plan:

Working in our Martian projects.

Anomalies in work:

None

Weather: Temperature: 2°C, Fairweather, Winds speed: NNE 22.5 MPH, Humidity 47%, Barometer 29.81 in

Crew Physical Status: Healthy

Today’s EVA:

Location: Candor Chasma

Duration: From 10:35 until 14:10

Purpose: Testing of long-range EVA protocols and testing of temperature regulation vest.

Today’s Extraordinary EVA:

Location: Intersection between White Moon and Factory Butte View

Duration: From 15:55 until 17:00

Purpose: Usage of rovers for battery discharge.

Reports to be filed: Sol Summary, GeenHab Report, Operations Report, EVA Request, EVA Report and Journalist Report.

Support Requested:

None

Greenhab Report – January 25th

GreenHab Report

Hernán David Mateus Jiménez

25/01/2018

Environmental control:

Heating

Ambient with door opening

Shade cloth on

Working Hour: 18:00
Inside temp at working hour: 14° C
Outside temp during working hours: 5° C
Inside temperature high: 30° C
Inside temperature low: 14° C
Inside humidity: 88 %RH

Inside humidity high: 88 %RH
Inside humidity low: 31 %RH

Hours of supplemental light:

For the crops 05:00 to 11:59 PM

Changes to crops: none

Daily water usage for crops: 8 gallons

Time(s) of watering for crops: 19:15

Changes to research plants: The plans of quinoa over the lysimeter are withered

Aquaponics: Atila is assembling it. He is waiting for the amount of water to fill the tank, this is necessary to do the test.

Narrative: none

EVA Report – January 23rd

EVA Report for 23Jan2018

Author: Oscar Ojeda

Purpose of EVA: Aerial mapping of terrain, testing of long-range EVA protocols, and long-range operation test of Cóndor Space Suit Simulator, Collection of possible halophile samples.

Participants: Oscar, Cynthia, Tania, and Danton

Coordinates: 518500 E, 4256500 N

Narrative: Two EVAs were undertaken today, the second one consisted on driving through Cow Dung Road to discharge the battery od each of the new rovers, to increase their battery life. The crew departed at 16:00 and was returning at the hab at 17:00. One of the crewmember’s radio was drained during the return, so he took the center position in the formation, for safety.

The first EVA took a crew of 4 to Lith canyon. The goals of the EVA were attained. The crew left the airlock at 10:35, securing the elements to the ATVs and taking the specified route, taking Cow Dung road north. The crew arrived at the entrance of the Special Region and struggled for about half an hour to find the entrance to the rest of the road, which is not easily visible. After finding the entrance the crew proceeded with the rest of the vehicle road, taking Lith Canyon Road to it’s end. At this point, the ATVs were left and the crew started walking up a hill to have a better visual of the area and plan the next exploration steps. Aerial footage was also taken from the Drone. Walking exploration of Lith canyon was undertaken, walking west from the position where the ATVs were parked. The crew walked towards the western wall of the canyon, departing after into the eastern limit, where aerial footage was gathered. The zone presented a significant amount of cactus and hervibore presence evidence. The crew started the return walk by 14:05, arriving to the ATVs by 14:45. The ATVs were boarded, and the road back resumed. On the way back, a sample of soil containing salt deposits was taken for further analysis for Halophiles. The crew arrived to the hab at 15:35, ending the EVA.